OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 2x + 5y = 4? A. (3, –2) B. (7, –2) C. (0, 1) D. (0, –1)

OpenStudy (anonymous):

2x+5y=4 (3,-2) 2(3)+5(-2)=4 6-10=4 -4=4 (7,-2) 2(7)+5(-2)=4 14-10=4 4=4 (0,1) 2(1)+5(0)=4 2+0=4 2=4 (0,-1) 2(-1)+5(0)=4 -2+0=4 -2=4

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

yep